OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for salsa classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 10(1.18)x The average rate of change in the number of students who enrolled for salsa classes from the second to the fifth year is ________students per year. Round your answer to the nearest whole number. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@SolomonZelman ,@Nnesha

OpenStudy (anonymous):

I want to make sure I have it right. I got 2.99 for my final answer but I rounded up and got 3. Is that correct?

OpenStudy (thehumantorch10):

In the second year, no of students who enrolled class is f(2) = 10 * 1.18^2 = 13.924 ~ 14 (students) Similarly, in fifth year, f(5) = 10 * 1.18^5 = 22.87 ~ 23 (students) The average rate of change from 2nd to 5th year is f(5)−f(2)5−2=3 Average rate of change is 3 students per year

OpenStudy (thehumantorch10):

you are correct

OpenStudy (anonymous):

Thank you so much. I just wanted to make sure and you made it very clear. Thank you.